OpenStudy (anonymous):

What belongs in the box marked with the question mark in the proof? Given: 4x + 1 = 6x - 2, Prove: x = 3/2 Statement Reason (1) 4x + 1 = 6x - 2 (2) 4x + 3 = 6x (3) 3 = 2x (4) 3/2 = x ? (5) x = 3/2

OpenStudy (anonymous):

\[1+4 \frac{3}{2}=-2+6 \frac{3}{2} \]7=7